OpenStudy (anonymous):

What is the equation of a line that is perpendicular to the line y=1/2x-8 and passes through the point (1, -6)? A.y=-2x-4 B.y=-1/2x-11/2 C.y=2x-8 D.y=2x+4

OpenStudy (loser66):

what is the slope of the given line?

OpenStudy (anonymous):

thats all it said

OpenStudy (loser66):

I know, I guide you how to find the answer. The very first step is to find out the slope of the given line y = 1/2 x-8. What is it? Hint: the number in the front of x

OpenStudy (anonymous):

1/2

OpenStudy (loser66):

good, so, the perpendicular line has the slope which is inverse of that number I mean, the slope of the required line is \(\dfrac{-1}{1/2}=?\)

OpenStudy (anonymous):

do what now

OpenStudy (loser66):

Wow, simplify the fraction, what do you have?

OpenStudy (anonymous):

bro ik im an idot thats y im on here idk how to do this and i got -2

OpenStudy (loser66):

yes, you are correct. And then, how many options have -2 in the front of x?

OpenStudy (anonymous):

only on yay thanks

OpenStudy (loser66):

got it? good